can someone please help me, lindsey has $6.25 in nickels and pennies for a total of 100 coins. How many of each? Please help me; thank you.
Rudy

Answer by Nate(3500)

LET US DEFINE VARIALBES:
n = amount of nickels
p = amount of pennies
--------------------------------------------------------------------
LET US COMPOSE EQUATIONS:
.05n + .01p = 6.25
n + p = 100
--------------------------------------------------------------------
LET US COMBINE THE EQUATIONS:
(.05n + .01p = 6.25)100
5n + p = 625
p = 625 - 5n
.............................
n + p = 100
n + 625 -5n = 100
-4n = -525
n = (525/4)
DOESN'T SEEM RIGHT......
I DON' THINK THERE IS A COMBINATION